OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (youngstudier):

Okay, what do you think you should first do? :)

OpenStudy (anonymous):

Subtract 16

OpenStudy (campbell_st):

this problem has no solution subtract 16 from both sides of the equation and you get \[\left| x + 2 \right| = -2\] doesn't work

OpenStudy (youngstudier):

Yep, from both sides so: 14-16 = -2 Then what do you think you should do?

OpenStudy (anonymous):

That's what I was thinking. It had no solution. And I would honestly check to see if it's right.

OpenStudy (youngstudier):

No wait, there is a solution.

OpenStudy (anonymous):

How?

OpenStudy (youngstudier):

Subtract 2 from both sides. -2 - 2 = -2 + (-2) = -4 And since it's the absolute value of i, then it would be positive 4. i = 4

OpenStudy (youngstudier):

Does that make sense?

OpenStudy (anonymous):

No not really... :/

OpenStudy (youngstudier):

Uhmmm.... l x + 2 l + 16 = 14 l x + 2 l + 16 - 16 = 14 -16 l x + 2 l = -2 l x + 2 - 2 l = -2 - 2 l x l = -4 x = 4

OpenStudy (youngstudier):

It's the absolute value of x, which makes the answer positive.

OpenStudy (campbell_st):

well I think @YoungStudier is attempting to use the identity that \[i^2 = -1\] it's used in complex numbers so you would have \[\left| x + 2 \right| = 2i^2\] so then you get 2 solutions \[x = -2 \pm 2i^2\] but I'd expect that you are working with real numbers

OpenStudy (youngstudier):

Uhm..... No, well, I don't know, I'm just solving it the way K12 taught me to. :P

OpenStudy (campbell_st):

no, you are asked to find a value for x, that when added to 2 and the absolute value is taken, the asnwer is -2 it doesn't exist

OpenStudy (campbell_st):

well is someone taught you to solve it in that fashion, I'd suggest you ask to change classes...

OpenStudy (youngstudier):

@Yungwisdom2000 , are you doing K12?

OpenStudy (campbell_st):

and when you solve an absolution value equation such as \[\left| x - 5 \right| = 10\] there are 2equations to solve 1. x - 5 = 10 and 2. x - 5 = -10 both solutions are valid.

OpenStudy (anonymous):

K12?? No. And I'm lost right about now